# Beyond Gaze Points: Augmenting Eye Movement with Brainwave Data for Multimodal User Authentication in Extended Reality

### v1

@@ 5.2. Overall Results
False Acceptance Rate (FAR) Insights: In practical applications, EER is often not the main metric of focus. Instead, the practical usage needs to minimize FAR to bolster system security against zero-effort attacks while maintaining a reasonable FRR To elucidate the trade-offs between these metrics, we present data in Table 2. A comparison between FAR at 0.0%, 0.1%, and 1% reveals that single-biometric approaches suffer from high FRRs, particularly in stringent security settings with low FARs (0.1% and 0.0%). For instance, in the S3 scenario with a FAR of 0.0%, FRRs for single-biometric approaches like Brainwaves and Eye-tracking with pupil are 67.99% and 25.00%, respectively. These high FRRs indicate that single-biometric methods could be impractical for high-security applications. In contrast, multi-modal methods, especially those utilizing score fusion, significantly alleviate this issue. For example, the Mean Score Fusion with and without pupil diameter method results in a much lower FRR of 3.789% and 11.61% at a 0.0% FAR, showcasing its effectiveness in balancing security and usability.
@@
In summary, our choice for both eye movement and brainwave data resulted in a significant improvement in the performance of our authentication system. Thus, we see our hypothesis confirmed that multimodal authentication based on eye movement and brainwaves substantially enhances the reliability of authentication using consumer-grade equipment with low sample recorder rates.
@@ 6.1. Technical Feasibility of XR Integration
For eye-tracking, we used the Pupil Core device with a 200Hz sample rate in our experiments. Prior research by Pastel et al. (Pastel et al., 2023) demonstrates the viability of integrating such technology in XR, citing 38 papers that used eye-trackers integrated with head-mounted displays (HMD). Moreover, companies like Pupil Labs offer eye-tracking solutions designed for AR and VR devices. Additionally, Varjo provides a VR headset with an integrated eye-tracker operating at a 200Hz sample rate, aligning with our hardware specifications [^6] https://varjo.com/products/vr-3/.
Similarly, for brainwave recording, we employed the Emotiv EPOC X, a consumer-grade device with a sampling rate of 256 Hz. Recently, Li et al. (Li et al., 2023) demonstrated that EEG sponge electrodes can be seamlessly integrated into VR headsets. Additionally, dry EEG electrodes have been developed and are commercially available [^7] https://www.bitbrain.com/neurotechnology-products/dry-eeg [^8] https://www.neurospec.com/Products/Details/1078/dsi-7 and some of these electrodes are already incorporated into VR devices [^9] https://www.neurospec.com/Products/Details/1077/dsi-vr300. Soon there will be VR headsets available that incorporate both EEG and eyetracking capabilities [^10] https://galea.co/.
@@ 9. Conclusion and Future Work
In this study, our investigation shows that this combination of brainwaves and eye movement yields highly promising results. Through our research, we substantially improved authentication accuracy and enhanced resistance against zero-effort attacks. Specifically, our multimodal authentication system achieved an EER of 0.298% and 0.686%, along with FRR of 3.8% and 11.6% at FAR of 0% compared with 25% and 52.8% FRR in single modality eye movement authentication. These results offer a higher level of security with a reasonable FRR, ensuring a smooth user experience without unnecessary disruptions. We provide a straightforward, hands-free authentication method that is both suitable for XR settings and appropriate for consumer-grade devices. Our multi-modal authentication system improves authentication accuracy while also holding promise for broader adoption in real-world applications.
In the future, investigating multi-session scenarios and increasing the sample size are essential steps for improving model learning and conducting more comprehensive evaluations. Furthermore, exploring additional tasks alongside the interactive dot task will contribute to a more extensive understanding of multimodal biometric authentication.

### v2

@@ 5.3. Overall Results
False Acceptance Rate (FAR) Insights: In practical applications, EER is often not the main metric of focus. Instead, the practical usage needs to minimize FAR to bolster system security against zero-effort attacks while maintaining a reasonable FRR To elucidate the trade-offs between these metrics, we present data in Table 2. A comparison between FAR at 0.01%, 0.1%, and 1% reveals that single-biometric approaches suffer from high FRRs, particularly in stringent security settings with low FARs (0.1% and 0.01%). For instance, in the S3 scenario with a FAR of 0.01%, FRRs for single-biometric approaches like Brainwaves and Eye-tracking with pupil are 67% and 25%, respectively. These high FRRs indicate that single-biometric methods could be impractical for high-security applications. In contrast, multi-modal methods, especially those utilizing score fusion, significantly alleviate this issue. For example, the Mean Score Fusion with and without pupil diameter method results in a much lower FRR of 3.789% and 11.61% at a 0.01% FAR, showcasing its effectiveness in balancing security and usability.
@@
Figure 5. Density distribution of brainwave similarity scores for legitimate users, human attackers, and random input features (it is similar for eye movement and fusion).
@@
Random input attack The density distance plot (Figure 5) demonstrates that the distances between randomly generated features and legitimate users are significantly greater than those between human attackers and legitimate users. This observation confirms that distance-based recognition systems are robust against random feature input attacks (Pagnin et al., 2014; Zhao et al., 2020), suggesting that such attacks are less effective than blind attacker scenarios.
Correlation between modalities: to investigate the correlation between synchronized eye movements and brainwaves, we applied the Pearson correlation coefficient  (Cohen et al., 2009) for time domain analysis and magnitude-squared coherence, calculated using Welch’s method (Welch, 1967), for frequency domain analysis. We explored correlations between: (1) different channels of eye movement and brainwave data separately; (2) synchronized brainwave and eye movement data; (3) brainwave and eye movement data across different times of the experiment; and (4) brainwave and eye movement data from different subjects. The analysis revealed a consistent trend in both the time and frequency domains, demonstrating a strong [^10] Higher than 0.6 correlation within the eye movement data and within the brainwave data (1). However, cross-modality correlations between eye movements and brainwaves were generally weak [^11] Lower than 0.1. Notably, the correlation of synchronized data from the same subject (2) was approximately 1.5 times stronger than that observed in cross-subject and time comparisons (3-4). The synchronized correlations demonstrated a more significant association between x-axis eye movement features and the frontal lobe region of the brain, particularly in the right hemisphere.
In summary, the weak correlation in synchronized data and strong correlation within modalities explain why the fusion of brainwave and eye movement data significantly improves the performance of our authentication system. Thus, we see our hypothesis confirmed that multimodal authentication based on eye movement and brainwaves substantially enhances the reliability of authentication using consumer-grade equipment with low sample recorder rates.
@@ 6.1. Technical Feasibility of XR Integration
For eye-tracking, we used the Pupil Core device with a 200Hz sample rate in our experiments. Prior research by Pastel et al. (Pastel et al., 2023) demonstrates the viability of integrating such technology in XR, citing 38 papers that used eye-trackers integrated with head-mounted displays (HMD). Moreover, companies like Pupil Labs and Tobii offer eye-tracking solutions designed for AR and VR devices  [^12] https://www.tobii.com/products/integration/xr-headsets. Additionally, Varjo provides a VR headset with an integrated eye-tracker operating at a 200Hz sample rate, aligning with our hardware specifications [^13] https://varjo.com/products/vr-3/.
Similarly, for brainwave recording, we employed the Emotiv EPOC X, a consumer-grade device with a sampling rate of 256 Hz. Recently, Li et al. (Li et al., 2023) demonstrated that EEG sponge electrodes can be seamlessly integrated into VR headsets. Additionally, dry EEG electrodes have been developed and are commercially available [^14] https://www.bitbrain.com/neurotechnology-products/dry-eeg [^15] https://www.neurospec.com/Products/Details/1078/dsi-7 and some of these electrodes are already incorporated into VR devices [^16] https://www.neurospec.com/Products/Details/1077/dsi-vr300. Soon there will be VR headsets available that incorporate both EEG and eye-tracking capabilities [^17] https://galea.co/.
@@ 7.3. Multimodal authentication
7.4. Comparison with VR Authentication Works
Reliable authentication in XR environment remains an open challenge. Biometrics have emerged as a viable solution; however, certain biometric methods require physical activities that may reduce practicality, ranging from discrete hand gestures (Peng et al., 2017; Chauhan et al., 2014) to more active ones like walking (Shen et al., 2019; Pfeuffer et al., 2019) or throwing a virtual ball (Ajit et al., 2019; Miller et al., 2019). Additionally, some solutions utilize sensors that may not be well-suited for XR environments (Chen et al., 2021; Gao et al., 2019); for instance, Chen et al. (Chen et al., 2021) employed electrical muscle stimulation, which relies on sensors attached to the hands.
While several studies propose biometrics suitable for XR setups, such as free head and body movement (Mustafa et al., 2018), eye-related biometrics (Zhang et al., 2018; Luo et al., 2020), skull conductance (Schneegass et al., 2016), and brainwaves (Lin et al., 2018b); their reported EERs range from 2.5% to 7%, comparable to our single-modality results (1.9%-4.9%). However, as shown in Table 2 , higher security configurations result in increased FRR. Therefore, the multimodal approach can be a promising alternative to improve performance further.
Multi-factor approaches (Zhu et al., 2020; Lu et al., 2018), such as combining biometrics with knowledge-based methods, aim to address performance challenges but inherit the limitations of both factors. An alternative is multimodal biometrics (Peng et al., 2016; Pfeuffer et al., 2019), which uses multi biometrics. We adopted this approach by fusing eye movement and brainwaves, reducing the error significantly.
@@ 9. Conclusion and Future Work
In this study, our investigation shows that this combination of brainwaves and eye movement yields highly promising results. Through our research, we substantially improved authentication accuracy and enhanced resistance against zero-effort attacks. Specifically, our multimodal authentication system achieved an EER of 0.298% and 0.686%, along with FRR of 3.8% and 11.6% at FAR of 0.01% compared with 25% and 52.8% FRR in single modality eye movement authentication. These results offer a higher level of security with a reasonable FRR, ensuring a smooth user experience without unnecessary disruptions. We provide a straightforward, hands-free authentication method that is both suitable for XR settings and appropriate for consumer-grade devices. Our multi-modal authentication system improves authentication accuracy while also holding promise for broader adoption in real-world applications.
In the future, investigating multi-session scenarios and increasing the sample size are essential steps for improving model learning and conducting more comprehensive evaluations. Furthermore, exploring additional tasks alongside the interactive dot task will contribute to a more extensive understanding of multimodal biometric authentication. Moreover, While the dot task can effectively resist against replay attacks, it would be beneficial to explore whether and how it is possible to use correlation in synchronized data to ensure that brainwave and eye movement data are recorded simultaneously, adding an extra layer of security.